A particle is moving eastwards with a velocity of 5m/s.In 10sec the velocity changes to 5m/s northwards.The average acceleration inthis time is

Grade:12

4 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

Vi =5i

Vf=5j

Avg acceleration =|(Vf -Vi)|/10

                        =|5j-5i)|10

                        =1/√2 m/s2


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

ASh
6 years ago
Dear pallavi 


Vi =5i

Vf=5j 

Avg acceleration =|(Vf -Vi)|/10

=|5j-5i)|10=1/√2 m/s2

 

SEKHAR
5 years ago
initial velocity =u=5m/s towards East 
Final velocity =v=5m/s towards North
Resultant velocity =R=dv=(52+52)1/2=7.071
Time =dt=10 sec
Now acceleration =dv/dt=7.071/10=0.7071
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your problem.
 
Vi =5i

Vf=5j

Avg acceleration = |(Vf -Vi)|/10

                        = |5j-5i)|10

                        = 1/√2 m/s2

Hope it helps.

Thanks and regards,

Kushagra